OpenStudy (anonymous):

If f(x) = log2 (x + 4), what is f−1(3)? @myininaya

OpenStudy (anonymous):

set \[\log_2(x+4)=3\] and solve for \(x\) first write in equivalent exponential form as \[2^3=x+4\] then solve for \(x\) in one step

OpenStudy (solomonzelman):

\[Log_2(x+4)=Log_2(2^3)\]

OpenStudy (solomonzelman):

Identities used.\[1=Log_cc\]\[3=3 \times Log_cc\]\[3=Log_c(c^3)\]